A University is interested in discovering what proportion of students are satisfied with the online delivery of modules. A total of 500 students were randomly selected from the academic administration register. Of these, 328 responded with 262 saying that they are satisfied with the online delivery. Produce a 95% confidence interval for the proportion of students who are satisfied.
Answer #1

Ans:

sample size,n=328 (as only 328 responded)

sample proportion=262/328=0.799

95% confidence interval for the proportion of students who are satisfied

=0.799+/-1.96*sqrt(0.799*(1-0.799)/328)

=0.799+/-0.043

=(0.755, 0.842)